Multiply 6/13 by the reciprocal of -7/16 
Multiply 6/13 by the reciprocal of -7/16 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Reciprocal of -7/16 = 16/-7 = -16/7 According to the question,
6/13 × (Reciprocal of -7/16)
6/13 × (-16/7) = -96/91

​Thanks